OpenStudy (anonymous):

evaluate log_ { 3 } 27 ^{9} =

OpenStudy (anonymous):

27^9=(3^3)^9=3^27, log_3(3^27)=27 Do you understand?

OpenStudy (anonymous):

Oh, come on, Luis, that is the opposite of elegant. :| Why, oh why, is there a massive number involved?